OpenStudy (anonymous):

f(x) = 3x + 1; g(x) = 5x - 1 Find f/g. (f/g)(x) = ; domain {x|x ≠ - } (f/g)(x) = ; domain {x|x ≠ - } (f/g)(x) = ; domain {x|x ≠ } (f/g)(x) = ; domain {x|x ≠ }

OpenStudy (nincompoop):

f/g =(3x+1)/(5x-1)

OpenStudy (nincompoop):

do you know what to do next?

OpenStudy (anonymous):

no:(

OpenStudy (agent0smith):

We already have f/g, now we need to find the values that aren't in the domain. You can't divide by zero, right? So what happens if the denominator of\[\Large f/g =\frac{3x+1} {5x-1}\] equals zero? Then we're dividing by zero... so we can't have 5x-1=0.

OpenStudy (anonymous):

ohh okay

OpenStudy (agent0smith):

So can you now find the excluded x value? (the one that makes 5x-1=0)

OpenStudy (anonymous):

yeah